What should be the next number in the following series? 100 96 104 88 120 56?

Answer by Cromlix(4381)

Answer is 184.
100 - 96 = 4
96 - 104 = -8
104 - 88 = 16
88 - 120 = -32
120 - 56 = 64
56 - 184 = -128.
Hope this helps.
:-)